### v2

# Automatically Evaluating the Paper Reviewing Capability of Large Language Models

Abstract
Peer review is essential for scientific progress, but it faces challenges such as reviewer shortages and growing workloads. Although Large Language Models (LLMs) show potential for providing assistance, research has reported significant limitations in the reviews they generate. While the insights are valuable, conducting the analysis is challenging due to the considerable time and effort required, especially given the rapid pace of LLM developments. To address the challenge, we developed an automatic evaluation pipeline to assess the LLMs’ paper review capability by comparing them with expert-generated reviews. By constructing a dataset [^1] https://figshare.com/s/d5adf26c802527dd0f62 consisting of 676 OpenReview papers, we examined the agreement between LLMs and experts in their strength and weakness identifications. The results showed that LLMs lack balanced perspectives, significantly overlook novelty assessment when criticizing, and produce poor acceptance decisions. Our automated pipeline enables a scalable evaluation of LLMs’ paper review capability over time.
Automatically Evaluating the Paper Reviewing Capability of
Large Language Models
Hyungyu Shin, Jingyu Tang, Yoonjoo Lee, Nayoung Kim, Hyunseung Lim, Ji Yong Cho§, Hwajung Hong, Moontae Lee§,∥, Juho Kim KAIST  Huazhong University of Science and Technology §LG AI Research  University of Illinois Chicago
1 Introduction
Figure 1: We consider paper review task that generates a summary of paper, strengths, weaknesses, and the final judgement. Our goal is to examine the level of agreement between LLMs and human experts in reviewing papers, based on their feedback targets and aspects.
Reviewing academic papers lies at the heart of scientific advancement, but it demands substantial expertise, time, and effort. The peer review system faces several challenges, including a growing number of submissions that outpace the reviewer availability, lack of incentives, and reviewer fatigue (Tropini et al., 2023; Horta and Jung, 2024; Hossain et al., 2025). While Large Language Models (LLMs) hold the potential to assist reviewers by reviewing papers automatically (Hosseini and Horbach, 2023; Robertson, 2023), prior research has reported significant limitations in their performance (Du et al., 2024; Liang et al., 2024; Zhou et al., 2024). For example, studies have highlighted that LLM-generated reviews often lack accuracy, detail, and specificity when compared to reviews written by human experts (Zhou et al., 2024; Mostafapour et al., 2024).
While these findings are informative, they do not sufficiently clarify the precise differences between expert-generated reviews and LLM-generated reviews. Specifically, it is still unclear to what extent LLMs excel or fall short in different aspects of reviewing compared to human experts. Addressing the question requires systematic quantitative analysis of review data, but such analysis is challenging to scale due to the significant time and effort required from researchers.
To address this gap, we introduce an automated pipeline designed to systematically analyze these differences. Our approach is to automatically annotate the strengths and weaknesses identified in reviews based on targets (e.g., problem, methodology, and experiment) and their associated aspects (e.g., validity, clarity, and novelty) and examine the agreement between LLMs and human experts (Figure 1). By introducing a systematic framework for examining the strengths and limitations of LLMs in academic review, this work offers valuable insights into improving their performance and enhancing their potential role in assisting the review process.
Figure 2: The overall evaluation process. Given a paper, we extract strengths and weaknesses from review data on the OpenReview platform. To identify key strengths and weaknesses that influence the final acceptance, we extracted them from the meta-review and augmented details from reviewer comments to make them self-contained. We then compared these with LLM-identified strengths and weaknesses, based on their feedback target and aspect. The evaluation is conducted automatically, enabling a scalable and longitudinal evaluation over time.
Our study leverages a dataset of 676 papers and their review data that has been collected from OpenReview [^2] https://openreview.net/ for ICLR conferences spanning 2021 to 2024. We extracted the strengths and weaknesses highlighted in the meta-reviews using gpt-4o333gpt-4o-2024-08-06, which will be compared with LLM-identified strengths and weaknesses. Then, we developed a coding schema (Table 1) for annotating the targets and aspects of these strengths and weaknesses, by surveying 9 AI paper submission guidelines and prior research on review analysis (Chakraborty et al., 2020; Ghosal et al., 2022; Yuan et al., 2022). Based on the schema, we manually annotated 327 strengths and weaknesses of 68 randomly sampled papers, providing a basis for building an automatic annotation tool. Our LLM-powered automatic annotation tool achieved 0.85 (target) and 0.86 (aspect) inter-rater reliability (IRR) with human-annotated results, showing high level of consistency and accuracy. Overall, we identified 3,657 review items (1,231 strengths and 2,416 weaknesses) from the review dataset.
We evaluated 8 LLMs (4 GPT, 2 Llama, and 2 DeepSeek family) for their paper review capability. After generating reviews for each of the 676 papers, we analyzed the agreement between LLMs’ and experts’ reviews based on their target and aspect assigned to strengths and weaknesses. The results showed that: 1) LLMs lack balanced perspective compared to human experts, 2) LLMs significantly neglect novelty assessment for evaluating papers’ weaknesses, and 3) the paper acceptance decisions are not accurate. The findings are consistent for all the LLMs, highlighting clear opportunities for improving their reasoning capability.
We release a dataset comprising 68 papers, experts reviews, 3,657 strengths and weaknesses identified from the reviews with automatically annotated targets and aspects, LLM-generated reviews from 8 LLMs, and a total of 43,042 strengths and weaknesses identified from the LLMs with their annotated targets and aspects.
2 Task
We define the paper review generation task as follows: given a research paper, 1) summarize main points, 2) identify a list of strengths and weaknesses, and 3) predict the final acceptance of the paper. This task offers direct value to various user groups (e.g., authors who want to get initial feedback on their draft, or reviewers who want to examine diverse view points) by providing actionable feedback for improving their papers. While valuable, evaluating papers based on research standards (e.g., novelty, rigor, and clarity) is difficult for LLMs as it requires significant expertise.
3 Constructing Expert Review Dataset
3.1 Collecting Review Data
We used real-world review data covering ICLR 2021-2024 from the OpenReview platform [^4] The review data is publicly available and permits use of data for research., where human experts evaluated submissions for a top-tier AI conference. Using the OpenReview API [^5] https://docs.openreview.net/getting-started/using-the-api and the list of submissions from public GitHub repositories [^6] https://github.com/{evanzd/ICLR2021-OpenReviewData, fedebotu/ICLR2022-OpenReviewData, fedebotu/ICLR2023-OpenReviewData, hughplay/ICLR2024-OpenReviewData} , we initially collected 18,407 submissions with their review data.
3.2 Identifying Strengths and Weaknesses
One of the challenges in identifying the strengths and weaknesses of these papers is that each review consists of multiple blocks, including a meta-review and individual review texts from several reviewers. To address the challenge, our approach is to use meta-review, a final review from a qualified expert that summarizes reviews and highlights important strengths and weaknesses for supporting the final decision. As the meta-review does not capture all the details, we created self-contained strengths and weaknesses by 1) extracting them from the meta-review and 2) augmenting these extracted elements with detailed comments from individual reviews (non-meta). We designed a prompting chain that consists of three prompts (Appendix A.1.1). After excluding withdrawn submissions that lack meta-reviews, 14,922 submissions remained.
Target
Code Definition (The review addresses …)
Problem Motivation, task definitions, and problem statements.
Prior Research References and contextual positioning of the submission.
Method Proposed approach, techniques, algorithms, or datasets.
Theory Theoretical foundations, assumptions, proofs, or justifications.
Experiment Experimental setup, results, and analysis.
Conclusion Findings, implications, discussions, and takeaways.
Paper General targets of the paper without specifying a particular target
Aspect
Code Definition (The review addresses …)
Impact Significance or practical influence of the work.
Novelty Originality of the submission compared to prior research.
Clarity Readability, ambiguity, or communication aspects.
Validity Soundness, completeness, and rigor.
Not-specific Multiple targets without emphasis on a particular aspect.
Table 1: The coding schema. To identify codes for targets (i.e., what the review praises or critiques) and aspects (i.e., the specific elements of the target being evaluated), we surveyed 9 AI paper submission guidelines (Appendix  A.2.1) and prior research on review analysis (Chakraborty et al., 2020; Ghosal et al., 2022; Yuan et al., 2022).
4 Building an Automatic Annotator Based on Target and Aspect
The central goal of this paper is to analyze where LLMs excel and fall short in reviewing papers, compared to human experts. To achieve the goal, we 1) annotate each of the strengths and weaknesses identified by LLMs and experts and 2) examine the agreement between them based on the annotation results. The analysis offers insights into the distinct contributions and limitations of LLMs in reviewing papers, informing strategies to foster more effective human-LLM collaboration in reviewing papers.
4.1 Developing a Coding Scheme
Our focus in classifying strengths and weaknesses lies in two key dimensions: targets (i.e., what the review praises or critiques) and aspects (i.e., the specific elements of the target being evaluated). To build an initial codebook, we surveyed 9 AI paper submission guidelines (Appendix A.2.1) and extracted target-aspect pairs from each statement in the guidelines (e.g., “The paper should state the full set of assumptions of all theoretical results if the paper includes theoretical results.” yields the target Theory and aspect Completeness). We also reviewed related work on the analysis of paper review data (Chakraborty et al., 2020; Ghosal et al., 2022; Yuan et al., 2022). After identifying 33 targets and 13 aspects, we merged similar items to create simple and distinct categories, resulting in 7 targets and 4 aspects. Table 1 shows our final coding scheme.
4.2 Building an Automatic Classifier Based on Human Annotation
Based on the coding scheme, we annotated targets and aspects of strengths and weaknesses to produce ground truth for developing an automatic annotator. We randomly sampled 68 papers from our review dataset, yielding 327 instances of strengths and weaknesses. Two authors annotated each instance together, resolving any conflicts. Most conflicts arose when an instance illustrated multiple points. For example, an instance such as “**Technically sound with a strong foundation**: The paper’s technical foundation is evident in its bi-level optimization framework, … Technical novelty also arises from using supermartingale constraints on the barrier function …” could correspond to both Validity and Novelty aspect. Two authors finalized the annotation through discussions, focusing on the main point or root cause of the issue. In the example, we annotated Validity, as the strength mainly praises the technical soundness, as shown in the header.
We then designed prompts to automatically annotate the instances, assigning a target and aspect label to each. Specifically, we designed four prompts where each corresponds to one of the four combinations of target/aspect and strength/weakness A.2.2. Table 2 shows the inter-rater reliability (IRR) between author annotations and LLM annotations. Classification using o3-mini achieved the IRR scores of 0.85 for targets and 0.86 for aspects. Given the high IRR and its relatively low computational cost, we used o3-mini for the automatic annotation of both target and aspect in the main evaluation. Moreover, an examination of the confusion matrix (Appendix A.2.3) suggests that the errors tend to occur in semantically related categories, indicating that the misclassifications are not arbitrary but rather reflect subtle ambiguities inherent in the data.
Model Target Aspect
gpt-4o-mini 0.75 0.80
gpt-4o 0.87 0.83
o3-mini 0.85 0.86
Table 2: Inter-Rater Reliability between annotations of authors and LLMs for the target and aspect.
5 Evaluation
Overall Strength Weakness
Model F1 Precision Recall F1 Precision Recall F1 Precision Recall
DeepSeek-R1 0.373 0.314 0.460 0.341 0.254 0.520 0.400 0.379 0.424
o1-mini 0.359 0.283 0.491 0.331 0.232 0.578 0.385 0.343 0.439
o1 0.355 0.300 0.436 0.318 0.234 0.495 0.388 0.377 0.400
DeepSeek-V3 0.351 0.300 0.421 0.330 0.246 0.501 0.368 0.362 0.374
Llama-405B 0.350 0.323 0.381 0.349 0.279 0.465 0.350 0.371 0.331
gpt-4o 0.349 0.287 0.442 0.342 0.252 0.533 0.354 0.325 0.388
gpt-4o-mini 0.344 0.289 0.427 0.335 0.246 0.522 0.353 0.337 0.369
Llama-70B 0.339 0.302 0.388 0.338 0.260 0.481 0.341 0.350 0.332
Table 3: Overall performance of alignments on strengths and weaknesses between experts-identified and LLM-identified reviews. The metrics were computed by comparing the (target, aspect) set between experts’ and LLMs’ review. DeepSeek-R1 achieved the best performance, o1-mini achieved superior recall, and Llama-405B achieved superior precision, compared to other models.
The goal of our evaluation is to analyze the strengths and weaknesses of a given paper as identified by LLMs, comparing them with those identified by human experts. Note that our evaluation does not consider the correctness of the identified strengths and weaknesses because our focus is comparing perspectives in reviewing papers for both groups, not the content itself.
The evaluation is based on paper-review pairs. However, we excluded accepted submissions in the evaluation because OpenReview provides the camera-ready versions (post-review) rather than the submitted versions (pre-review), leading to a mismatch between the collected review and the camera-ready paper. Therefore, we only focused on rejected papers, where the meta-review corresponds to the latest version of the paper. Out of 9,139 rejected papers, we randomly sampled 7.5% of them (685 papers) for the evaluation. In total, we obtained 3,689 review items (1,241 strengths and 2,448 weaknesses), each automatically annotated with a target and aspect label.
Figure 3: Distribution of strengths and weaknesses. Unlike human experts, LLMs reported a consistent count regardless of paper contents. o1-mini identified the most, while Llama models identified the fewest points.
5.1 Large Language Models
We consider eight off-the-shelf LLMs, differing in size and availability (open-source vs. proprietary): four GPT models (gpt-4o-mini, gpt-4o, o1-mini, o3-mini, o1)777gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, o1-mini-2024-09-12, o1-2024-12-17, two Llama models (Llama 3.1-{70B, 405B}), and two DeepSeek models (DeepSeek-{V3, R1}). We used the default parameters of the models.
5.2 Procedure
For each of the 685 papers, we generated review data using the prompting pipeline (Section 3), extracted the strengths and weaknesses, and annotated the corresponding targets and aspects using the automatic annotator powered by o3-mini. Then for each LLM in Section 5.1, we generated a review for each paper (See Appendix A.1.2 for the prompt), extracted the strengths and weaknesses, and annotated the targets and aspects using the same automatic annotator. Then we compared the annotated targets and aspects between the experts’ reviews and LLMs’ reviews.
5.3 Result
Figure 4: Normalized distributions by target/aspect and strength/weakness for LLMs and human experts (red line). Overall, both groups showed similar perspectives in reviewing papers, focusing on technical targets (i.e., Method, Experiment, and Theory) and validity. However, LLMs showed more biased perspectives that focus on the technical validity whereas human experts exhibited more balanced perspectives. However, all the LLMs lack consideration of Novelty for weaknesses compared to human experts, which is a significant limitation in reviewing papers.
While human experts raised various number of points, LLMs identified a relatively consistent number of points regardless of the paper’s content. Moreover, LLMs identified a similar number of points between strengths and weaknesses, which was a different pattern from that of the human experts. Figure 3 shows the distribution of the number of strengths and weaknesses identified by human experts and LLMs. Overall, LLMs identified more points on average (7.88) than human experts (5.39). Among the LLMs, Llama models identified fewer (3.17 strengths and 3.15 weaknesses, on average) whereas o1-mini reported more strengths and weaknesses (5.03 and 5.47, respectively) than other models. By comparing target and aspect labels between human experts and LLMs, we report the following key findings.
Overall, LLMs do not effectively identify key targets and aspects when reviewing papers. Table 3 shows the overall performance of LLMs, which computes the agreement of (target, aspect) labels between human experts and LLMs. The best F1 score among the LLMs was 0.37, which indicates a low level of agreement with human experts in identifying strengths and weaknesses. Since we only considered whether the categories of review items match rather than their detailed content, the result implies that the actual content of strengths and weaknesses would be significantly different between human experts and LLMs. In general, LLMs showed higher recall than precision scores, mainly due to the nature of identifying a higher number of review points than human experts. Also, LLMs consistently achieved higher F1 scores for weaknesses than strengths. Among the LLMs, deepseek-r1 achieved the best overall performance, o1-mini achieved the best recall, and Llama-405B achieved the best precision.
While overall agreement is low, both groups primarily emphasized technical validity and novelty in the strengths, and focused on technical validity and clarity in the weaknesses. Figure 4 shows the normalized distribution of target and aspect labels for both experts and LLMs. For targets, both groups primarily focused on core technical elements—Method, Experiment, and Theory. However, strengths and weaknesses illustrated different patterns: both groups praised Method more than Experiment in the strengths, but criticized Experiment more than Method in the weaknesses. For aspects, both groups considered Validity as an important aspect, especially when evaluating weaknesses. Impact received more attention than Clarity in the strengths, whereas the opposite was observed in the weaknesses.
LLMs consistently exhibited a more biased perspective, while human experts maintained a more balanced perspective. Although both groups shared a core perspective, LLMs tend to focus on a few specific dimensions. For instance, LLMs focused primarily on Method and Experiment, while neglecting Prior Research (e.g., whether the paper adequately addresses prior work in positioning) and Problem (e.g., whether the task needs community attention), which human experts point out (Problem in the strengths and Prior Research in the weaknesses). For aspects, LLMs mostly focused on Validity in both strengths and weaknesses. In contrast, human experts considered the aspects more evenly among Validity, Novelty, and Clarity. Notably, LLMs exhibited a significant bias for Novelty. LLMs praised Novelty in the strengths, whereas they rarely criticized it in the weaknesses. This is a significant drawback, as a paper review requires a critical examination of novelty, by comparing them against existing work.
Table 4: F1 Score for target and aspects between DeepSeek-R1, o1-mini, and Llama-405B. Due to the biased perspective of LLMs, we observed a clear gap between what LLMs mostly focus on (e.g., Method and Experiment targets and Validity aspect) and overlook (e.g., Problem target and Novelty aspect). Full results (F1 score, precision, and recall across models and target/aspect labels) are available in Appendix A.2.4.
Target F1 score
Target DeepSeek-R1 o1-mini Llama-405B Problem 0.30 0.40 / 0.20 0.28 0.35 / 0.20 0.16 0.16 / 0.16   Method 0.73 0.75 / 0.71 0.76 0.75 / 0.77 0.69 0.76 / 0.63   Theory 0.47 0.44 / 0.51 0.47 0.41 / 0.53 0.43 0.46 / 0.40   Experiment 0.68 0.51 / 0.85 0.68 0.51 / 0.86 0.66 0.52 / 0.81
Aspect F1 score
Aspect DeepSeek-R1 o1-mini Llama-405B Novelty 0.39 0.66 / 0.12 0.39 0.66 / 0.12 0.34 0.66 / 0.01   Impact 0.41 0.54 / 0.29 0.43 0.56 / 0.30 0.32 0.35 / 0.29   Validity 0.77 0.60 / 0.95 0.77 0.60 / 0.95 0.77 0.60 / 0.95   Clarity 0.27 0.17 / 0.36 0.40 0.30 / 0.50 0.28 0.16 / 0.40
Due to their biased focus, the level of agreement between LLMs and human experts varied across different labels. Table 4 shows F1 scores for specific targets and aspects. For targets and aspects that LLMs focus more on — Method and Experiment targets and Validity aspect — LLMs had a much higher level of agreement with human experts compared to other targets and aspects. In the case of Experiment, the F1 score was consistently higher for weaknesses than strengths, suggesting that LLMs are more effective at identifying concerns in experiments (e.g., lack of baselines or scope of evaluation) than recognizing strong points of theories (e.g., experiments are rigorous and thorough). Similarly, for aspects other than Validity, agreement levels were notably lower. In particular, Novelty in the weaknesses, which LLMs largely overlooked, showed a significantly lower F1 score.
LLMs showed similar patterns in their emphasis in reviewing papers, regardless of their size and reasoning capability. All LLMs, including both proprietary and open source models, showed similar patterns that focused primarily on technical (Method, Experiment, and Theory) validity than on Novelty for the weaknesses. This consistency indicates that the observed biases could stem from the inherent design and training methods of LLMs, revealing potential room for improvement in the reasoning capability that requires leveraging external information (e.g., identifying comparable related work and analyzing novelty of submissions).
Final acceptance decisions are not accurate. Table 5 shows the rejection rate reported by each LLM. Overall, the best achieved rejection rate was 24.9%, which indicates that recall for rejected papers is poor. gpt-4o, Llama-405B, and DeepSeek-R1 performed significantly better than other models. gpt-4o and Llama-405B showed relatively high rejection rates while their agreement with human experts on strengths and weaknesses was low (0.348 and 0.349). DeepSeek-R1 demonstrated a moderate rejection rate among the models, with a relatively higher agreement score on strengths and weaknesses (0.373).
Table 5: Rejection percentage by model. 100% is the highest score, as we considered rejected papers. All the models were highly positive about the paper acceptance, although the papers were rejected.
Model Rejection (%)
gpt-4o 27.92%
Llama-405B 24.30%
DeepSeek-r1 23.79%
gpt-4o-mini 9.50%
DeepSeek-v3 7.93%
o1-mini 5.45%
o1 3.36%
Llama-70B 0.74%
6 Discussion
In this paper, we found gaps in the way of reviewing papers between human experts and LLMs and reported several limitations of LLMs as an automated reviewer, using an automated pipeline. Based on the results, we discuss the following implications.
There exists significant room for improving alignments between human experts and LLMs in reviewing papers. Our results show that LLMs exhibit a more biased perspective, which mostly examines technical validity without contextual consideration, compared to human experts. To reduce the gap, fine-tuning models using our dataset could serve as a starting point. While our results revealed significant limitations of LLMs in reviewing papers, our focus was mostly on the target and aspect labels rather than comparing actual content. We suspect that a more significant gap lies in the actual content addressed in the review items, even if they share the same target and aspect labels. For instance, (Experiment, Validity) could point out either lack of necessary baselines or lack of ablation studies to justify authors’ arguments. Content-level investigations may reveal more limitations of LLMs, ultimately contributing to improving the reasoning capability of LLMs.
Research should investigate the task of assessing the novelty of academic papers. Our finding illustrated that all LLMs in our analysis significantly overlooked the novelty aspect when evaluating weaknesses of papers. Previous studies have indicated that language models’ ability to assess novelty is inferior to that of experts Julian Just and Hutter (2024); Lin et al. (2024), emphasizing the need to encourage LLMs to focus on novelty evaluation. Although novelty is one of the most important aspects in reviewing papers and efforts have been made to enhance LLMs’ ability to assess novelty (Bougie and Watanabe, 2024; Lin et al., 2024), there exists no suitable benchmark for systematically measuring the novelty assessment capability of LLMs. We believe that creating the benchmark is a valuable contribution to the field, which allows LLMs to learn how to assess similarities between papers. Leveraging data in OpenReview could be an initial step as it contains experts’ judgement on novelty of the paper for both positive and negative decisions.
It is important to further explore the alignment between points of strengths and weaknesses and the final decision based on them. We found that there exists a discrepancy between achieving a high level of agreement in strengths and weaknesses and correctly predicting the final decision. It implies that the way LLMs make the final decision based on the identified strengths and weaknesses could be different from the way of human experts. Consistent with our findings, previous research revealed that LLMs offer positive assessments Latona et al. (2024a). As it is important to inform clear and convincing rationale behind the final decision, further investigation is needed to carefully evaluate whether the final decision based on the strengths and weaknesses is reasonable, for various stakeholders such as domain experts or novices. Generating reviews with clear alignments between the review points and final decision is a challenging task because often the relationship is not very clear and implicit Zhou et al. (2024), relying on community norms and social factors. Learning the relationship from the review data could be useful for understanding the gap between human experts and LLMs in decision-making.
7 Related Work
With the powerful reasoning capability of LLMs, LLMs have the potential to assist in the task of reviewing papers (Latona et al., 2024b; D’Arcy et al., 2024). Research has explored the capability of LLMs in reviewing papers, identifying a set of limitations. While LLM-generated reviews can be helpful (Liang et al., 2024; Tyser et al., 2024; Lu et al., 2024), research has shown that LLMs-generated reviews lack diversity (Du et al., 2024; Liang et al., 2024) and technical details (Zhou et al., 2024), exhibit bias (Ye et al., 2024), tend to provide positive feedback (Zhou et al., 2024; Du et al., 2024), and may include irrelevant or even inaccurate comments (Mostafapour et al., 2024). Furthermore, research also has reported that LLM-generated reviews have a low level of agreement with experts-generated reviews (Saad et al., 2024).
To assess the quality of review, research has taken a quantitative approach by analyzing review text. For instance, research has evaluated the quality of review based on human preferences (Tyser et al., 2024), similarity to human-generated review (Zhou et al., 2024; Liang et al., 2024; Gao et al., 2024; Sun et al., 2024; Chamoun et al., 2024) and classification-based scores (Li et al., 2023). Another approach is to classify review data based on categories such as section (Ghosal et al., 2022), aspect (Yuan et al., 2022; Chamoun et al., 2024; Liang et al., 2024) and actionability (Choudhary et al., 2022). While quantitative approach provides concrete insights, it is typically conducted as a one-time evaluation, challenging to apply the consistent methodology over time.
8 Conclusion
We introduced an automatic evaluation pipeline to assess LLMs’ capability in reviewing papers, by examining the agreement between LLM-generated and expert-generated reviews based on their target and aspect annotations for strengths and weaknesses. Our findings suggest that LLMs need to adopt a more balanced perspective, place greater emphasis on novelty assessment when critiquing papers, and better formulate their final judgement based on the identified strengths and weaknesses. We believe that our automated pipeline can contribute to ongoing evaluation of LLMs’ paper review capabilities within the rapid pace of LLM developments, offering concrete insights for improving their reasoning capability.
Acknowledgments
This work was conducted with the support of LG AI Research, which is gratefully acknowledged.
Limitation
This paper has the following limitations. First, our dataset focuses solely on ICLR submissions and the coding schema is developed based on AI venues, which limit generalizability to other fields. Second, our analysis examines the target and aspect of the review items, but other important dimensions such as level of specificity and depth of justification remain unexplored. Third, while our automatic annotator achieved high IRR (0.85) with human annotations, some discrepancies still exist. Finally, we did not explore possible prompt engineering strategies that could mitigate the limitations of LLMs in paper review. Future work can investigate techniques to enhance the alignment between LLMs and human experts.
Ethical impact
This paper presents potential risks. First, while our vision is to build LLMs to effectively assist review process, our work could inadvertently encourage over-reliance on LLM-generated reviews among various user groups, including reviewers and novice researchers. Second, although our dataset could contribute to improving LLM performance of reviewing papers, it may introduce a certain bias due to the source of dataset; ICLR for papers and code based on AI research. Finally, we assess the quality of review based on alignment with expert reviews, but it could offer a potentially biased perspective, as our coding schema only considers two dimensions, which may undervalue the unique contributions of LLM-generated reviews.
Appendix A Appendix
A.1 Review Generation
A.1.1 Prompts for Expert Review Generation
In this section, we provide prompts for identifying key strength and weakness from review data. Figure 5 shows the prompt for extracting weakness and strength from meta-review. Figure 6 shows the prompt for using detailed comments from reviews to augment the extracted elements. Figure 7 shows the prompt for removing some extraneous reference. We used the three prompts in a prompt chain, sequentially running the prompts.
Figure 5: Prompt for Meta-Review Summarization
Figure 6: Prompt for Generating Augmented Review
Figure 7: Prompt for Paraphrasing Augmented Review
A.1.2 Prompts for LLM Review Generation
Figure 8 shows the prompt for using LLM to generate reviews from paper.
Figure 8: Prompt for LLM Review Generation
A.2 Details of Building Automatic Annotator
A.2.1 AI paper writing guidelines
To ensure guidelines are comprehensive, we collected guidelines from 9 sources, comprising a total of 243 items, as shown in Table 6. An item refers to a specific requirement mentioned in the guidelines, which serves as a distinct criterion for reviewing or writing a paper.
Table 6: Guidelines and Item Count Summary
Guideline Item Count
ICML Paper Writing Best Practices [^1] https://icml.cc/Conferences/2022/BestPractices 38
ICML 2023 Paper Guidelines [^2] https://icml.cc/Conferences/2023/PaperGuidelines 30
NIPS 2024 Reviewer Guidelines [^3] https://neurips.cc/Conferences/2024/ReviewerGuidelines 18
ACL Checklist [^4] https://aclrollingreview.org/responsibleNLPresearch/ 49
How to Write a Good Research Paper in the Machine Learning Area [^5] https://www.turing.com/kb/how-to-write-research-paper-in-machine-learning-area 6
ACL Ethics Review Questions [^6] https://2023.eacl.org/ethics/review-questions/ 21
AAAI Reproducibility Checklist [^7] https://aaai.org/conference/aaai/aaai-25/aaai-25-reproducibility-checklist/ 29
NeurIPS 2021 Paper Checklist Guidelines [^8] https://neurips.cc/Conferences/2021/PaperInformation/PaperChecklist 46
ICLR 2019 Guidelines [^9] https://iclr.cc/Conferences/2019/Reviewer_Guidelines 6
Total Count 243
A.2.2 Prompts
In this section, we provide prompts designed to annotate reviews. We designed 4 prompts where each corresponds to one of the four combinations of target/aspect and strength/weakness. Specifically, we designed Target-Strength (Figure 9), Aspect-Strength, (Figure 11), Target-Weakness (Figure 10) , and Aspect-Weakness (Figure 12) prompts.
Figure 9: Prompt for Automatic Target Annotation for Strength
Figure 10: Prompt for Automatic Target Annotation for Weakness
Figure 11: Prompt for Automatic Aspect Annotation for Strength
Figure 12: Prompt for Automatic Aspect Annotation for Weakness
A.2.3 Annotation Comparison
We present a comparison between LLM and human annotations for both target and aspect. Figures 13 and Figure 14 illustrate the discrepancies. Areas of alignment between LLM and human annotations are shown in green, while red highlights regions with significant discrepancies.
Figure 13: LLM vs. human target annotation
Figure 14: LLM vs. human aspect annotation
While LLM annotations differ from human annotations in some cases, certain discrepancies remain reasonable. Figure 15 and Figure 16 illustrate examples of such reasonable discrepancies.
Figure 15: Cases of Target Annotation Discrepancy
Figure 16: Cases of Aspect Annotation Discrepancy
A.2.4 Results
The following tables present a comprehensive performance comparison of models across different metrics and evaluation targets, including both strengths and weaknesses (Table 7), as well as separate analyses focusing on strengths (Table 8) and weaknesses (Table 9). Additionally, we provide a similar comparison across metrics and broader aspects, including both strengths and weaknesses (Table 10), strengths alone (Table 11), and weaknesses alone (Table 12).
Table 7: Performance Comparison of Models Across Metrics and Targets (Including both Strengths and Weaknesses)
Target Problem Prior Research Method Theory Experiment Conclusion Paper
F1 (gpt-4o-mini) 0.268 0.076 0.737 0.427 0.680 0.103 0.227
F1 (gpt-4o) 0.292 0.052 0.741 0.448 0.673 0.089 0.247
F1 (o1-mini) 0.275 0.054 0.764 0.472 0.684 0.175 0.253
F1 (o1) 0.274 0.044 0.754 0.489 0.673 0.133 0.091
F1 (llama-70B) 0.269 0.049 0.711 0.410 0.659 0.172 0.158
F1 (llama-405B) 0.158 0.031 0.690 0.427 0.662 0.167 0.134
F1 (deepseek-r1) 0.297 0.081 0.729 0.473 0.682 0.164 0.152
F1 (deepseek-v3) 0.241 0.051 0.725 0.405 0.680 0.110 0.092
Prec (gpt-4o-mini) 0.317 0.134 0.647 0.317 0.549 0.063 0.241
Prec (gpt-4o) 0.298 0.109 0.634 0.334 0.547 0.057 0.251
Prec (o1-mini) 0.315 0.130 0.639 0.342 0.549 0.107 0.274
Prec (o1) 0.279 0.064 0.648 0.381 0.549 0.111 0.245
Prec (llama-70B) 0.339 0.143 0.653 0.295 0.548 0.105 0.289
Prec (llama-405B) 0.324 0.071 0.647 0.310 0.558 0.115 0.233
Prec (deepseek-r1) 0.321 0.099 0.639 0.327 0.549 0.135 0.301
Prec (deepseek-v3) 0.288 0.100 0.645 0.280 0.547 0.076 0.249
Rec (gpt-4o-mini) 0.233 0.053 0.870 0.691 0.983 0.274 0.232
Rec (gpt-4o) 0.297 0.034 0.899 0.723 0.965 0.202 0.270
Rec (o1-mini) 0.266 0.034 0.952 0.834 0.994 0.536 0.249
Rec (o1) 0.353 0.034 0.905 0.736 0.963 0.167 0.056
Rec (llama-70B) 0.246 0.030 0.803 0.720 0.919 0.476 0.146
Rec (llama-405B) 0.108 0.020 0.774 0.694 0.894 0.300 0.095
Rec (deepseek-r1) 0.299 0.069 0.859 0.865 0.983 0.357 0.102
Rec (deepseek-v3) 0.210 0.035 0.844 0.755 0.981 0.238 0.058
Table 8: Performance Comparison of Models Across Metrics and Targets (Strengths)
Target Problem Prior Research Method Theory Experiment Conclusion Paper
F1 (gpt-4o-mini) 0.283 0.000 0.760 0.424 0.511 0.118 0.232
F1 (gpt-4o) 0.329 0.000 0.756 0.446 0.517 0.143 0.119
F1 (o1-mini) 0.345 0.000 0.753 0.411 0.511 0.300 0.233
F1 (o1) 0.384 0.000 0.749 0.470 0.512 0.267 0.061
F1 (llama-70B) 0.245 0.000 0.750 0.420 0.516 0.242 0.198
F1 (llama-405B) 0.160 0.000 0.755 0.455 0.516 0.333 0.079
F1 (deepseek-r1) 0.396 0.000 0.749 0.436 0.513 0.174 0.135
F1 (deepseek-v3) 0.331 0.000 0.755 0.423 0.509 0.114 0.086
Prec (gpt-4o-mini) 0.315 0.000 0.622 0.286 0.343 0.071 0.198
Prec (gpt-4o) 0.295 0.000 0.616 0.299 0.350 0.091 0.182
Prec (o1-mini) 0.314 0.000 0.611 0.264 0.343 0.176 0.203
Prec (o1) 0.285 0.000 0.624 0.322 0.346 0.222 0.172
Prec (llama-70B) 0.404 0.000 0.620 0.275 0.352 0.148 0.178
Prec (llama-405B) 0.419 0.000 0.620 0.319 0.358 0.231 0.163
Prec (deepseek-r1) 0.355 0.000 0.617 0.289 0.347 0.103 0.279
Prec (deepseek-v3) 0.364 0.000 0.620 0.276 0.344 0.069 0.154
Rec (gpt-4o-mini) 0.258 0.000 0.975 0.819 0.996 0.333 0.281
Rec (gpt-4o) 0.371 0.000 0.978 0.872 0.991 0.333 0.089
Rec (o1-mini) 0.382 0.000 0.980 0.935 0.996 1.000 0.274
Rec (o1) 0.588 0.000 0.936 0.872 0.987 0.333 0.037
Rec (llama-70B) 0.176 0.000 0.948 0.894 0.969 0.667 0.224
Rec (llama-405B) 0.099 0.000 0.965 0.796 0.921 0.600 0.052
Rec (deepseek-r1) 0.447 0.000 0.953 0.883 0.983 0.571 0.089
Rec (deepseek-v3) 0.303 0.000 0.963 0.904 0.982 0.333 0.059
Table 9: Performance Comparison of Models Across Metrics and Targets (Weaknesses)
Target Problem Prior Research Method Theory Experiment Conclusion Paper
F1 (gpt-4o-mini) 0.253 0.153 0.715 0.430 0.849 0.088 0.222
F1 (gpt-4o) 0.256 0.104 0.726 0.449 0.830 0.036 0.375
F1 (o1-mini) 0.204 0.108 0.774 0.534 0.857 0.050 0.272
F1 (o1) 0.164 0.089 0.760 0.508 0.835 0.000 0.120
F1 (llama-70B) 0.294 0.098 0.672 0.400 0.802 0.103 0.118
F1 (llama-405B) 0.155 0.062 0.625 0.399 0.809 0.000 0.190
F1 (deepseek-r1) 0.198 0.163 0.709 0.510 0.852 0.154 0.169
F1 (deepseek-v3) 0.151 0.103 0.696 0.387 0.850 0.105 0.099
Prec (gpt-4o-mini) 0.320 0.268 0.672 0.347 0.755 0.056 0.283
Prec (gpt-4o) 0.301 0.219 0.651 0.369 0.743 0.024 0.321
Prec (o1-mini) 0.315 0.259 0.666 0.420 0.754 0.038 0.345
Prec (o1) 0.273 0.127 0.672 0.440 0.752 0.000 0.317
Prec (llama-70B) 0.274 0.286 0.687 0.315 0.744 0.062 0.400
Prec (llama-405B) 0.228 0.143 0.673 0.300 0.758 0.000 0.304
Prec (deepseek-r1) 0.287 0.197 0.661 0.365 0.750 0.167 0.323
Prec (deepseek-v3) 0.212 0.200 0.669 0.284 0.750 0.083 0.345
Rec (gpt-4o-mini) 0.209 0.107 0.764 0.563 0.970 0.214 0.183
Rec (gpt-4o) 0.222 0.068 0.821 0.574 0.939 0.071 0.451
Rec (o1-mini) 0.151 0.068 0.924 0.732 0.992 0.071 0.224
Rec (o1) 0.118 0.068 0.874 0.600 0.939 0.000 0.074
Rec (llama-70B) 0.316 0.059 0.658 0.547 0.869 0.286 0.069
Rec (llama-405B) 0.118 0.040 0.583 0.593 0.867 0.000 0.138
Rec (deepseek-r1) 0.151 0.139 0.764 0.847 0.984 0.143 0.115
Rec (deepseek-v3) 0.118 0.069 0.725 0.605 0.980 0.143 0.057
Table 10: Performance Comparison of Models Across Metrics and Aspects (Including both Strengths and Weaknesses)
Aspect Novelty Impact Validity Clarity
F1 (gpt-4o-mini) 0.334 0.390 0.775 0.396
F1 (gpt-4o) 0.378 0.428 0.769 0.365
F1 (o1-mini) 0.386 0.427 0.773 0.395
F1 (o1) 0.404 0.399 0.772 0.401
F1 (llama-70B) 0.334 0.322 0.769 0.327
F1 (llama-405B) 0.337 0.318 0.772 0.278
F1 (deepseek-r1) 0.387 0.414 0.775 0.266
F1 (deepseek-v3) 0.346 0.422 0.768 0.187
Prec (gpt-4o-mini) 0.367 0.291 0.671 0.317
Prec (gpt-4o) 0.474 0.313 0.668 0.298
Prec (o1-mini) 0.528 0.300 0.668 0.311
Prec (o1) 0.589 0.305 0.669 0.334
Prec (llama-70B) 0.665 0.318 0.667 0.337
Prec (llama-405B) 0.587 0.302 0.671 0.332
Prec (deepseek-r1) 0.535 0.308 0.670 0.339
Prec (deepseek-v3) 0.504 0.306 0.664 0.309
Rec (gpt-4o-mini) 0.460 0.600 0.990 0.549
Rec (gpt-4o) 0.506 0.689 0.975 0.485
Rec (o1-mini) 0.507 0.758 0.990 0.548
Rec (o1) 0.435 0.579 0.981 0.511
Rec (llama-70B) 0.450 0.371 0.981 0.346
Rec (llama-405B) 0.478 0.352 0.978 0.241
Rec (deepseek-r1) 0.502 0.632 0.988 0.219
Rec (deepseek-v3) 0.478 0.683 0.982 0.134
Table 11: Performance Comparison of Models Across Metrics and Aspects (Strengths)
Aspect Novelty Impact Validity Clarity
F1 (gpt-4o-mini) 0.643 0.474 0.599 0.309
F1 (gpt-4o) 0.654 0.520 0.593 0.202
F1 (o1-mini) 0.656 0.556 0.592 0.299
F1 (o1) 0.626 0.530 0.596 0.342
F1 (llama-70B) 0.636 0.411 0.593 0.292
F1 (llama-405B) 0.660 0.345 0.596 0.157
F1 (deepseek-r1) 0.655 0.536 0.598 0.170
F1 (deepseek-v3) 0.660 0.547 0.585 0.122
Prec (gpt-4o-mini) 0.498 0.368 0.431 0.222
Prec (gpt-4o) 0.498 0.398 0.428 0.190
Prec (o1-mini) 0.501 0.403 0.424 0.224
Prec (o1) 0.530 0.412 0.430 0.261
Prec (llama-70B) 0.497 0.467 0.426 0.236
Prec (llama-405B) 0.506 0.368 0.431 0.215
Prec (deepseek-r1) 0.503 0.400 0.431 0.224
Prec (deepseek-v3) 0.509 0.403 0.419 0.207
Rec (gpt-4o-mini) 0.907 0.667 0.986 0.511
Rec (gpt-4o) 0.955 0.749 0.965 0.216
Rec (o1-mini) 0.949 0.897 0.979 0.449
Rec (o1) 0.763 0.744 0.969 0.496
Rec (llama-70B) 0.883 0.366 0.976 0.384
Rec (llama-405B) 0.949 0.324 0.969 0.123
Rec (deepseek-r1) 0.937 0.809 0.976 0.137
Rec (deepseek-v3) 0.940 0.851 0.965 0.086
Table 12: Performance Comparison of Models Across Metrics and Aspects (Weaknesses)
Aspect Novelty Impact Validity Clarity
F1 (gpt-4o-mini) 0.024 0.306 0.951 0.484
F1 (gpt-4o) 0.103 0.335 0.945 0.528
F1 (o1-mini) 0.116 0.299 0.954 0.492
F1 (o1) 0.182 0.268 0.949 0.459
F1 (llama-70B) 0.032 0.233 0.945 0.362
F1 (llama-405B) 0.013 0.291 0.947 0.399
F1 (deepseek-r1) 0.120 0.292 0.952 0.362
F1 (deepseek-v3) 0.031 0.297 0.951 0.253
Prec (gpt-4o-mini) 0.235 0.214 0.912 0.411
Prec (gpt-4o) 0.450 0.228 0.907 0.406
Prec (o1-mini) 0.556 0.197 0.911 0.397
Prec (o1) 0.647 0.198 0.908 0.406
Prec (llama-70B) 0.833 0.169 0.907 0.438
Prec (llama-405B) 0.667 0.236 0.911 0.450
Prec (deepseek-r1) 0.568 0.215 0.908 0.454
Prec (deepseek-v3) 0.500 0.209 0.908 0.410
Rec (gpt-4o-mini) 0.013 0.533 0.994 0.587
Rec (gpt-4o) 0.058 0.630 0.985 0.754
Rec (o1-mini) 0.065 0.619 1.000 0.646
Rec (o1) 0.106 0.415 0.994 0.527
Rec (llama-70B) 0.016 0.376 0.987 0.308
Rec (llama-405B) 0.006 0.381 0.987 0.359
Rec (deepseek-r1) 0.067 0.455 1.000 0.302
Rec (deepseek-v3) 0.016 0.515 0.998 0.183