### v3

# Mind the Blind Spots: A Focus-Level Evaluation Framework for LLM Reviews

Abstract
Peer review underpins scientific progress, but it is increasingly strained by reviewer shortages and growing workloads. Large Language Models (LLMs) can automatically draft reviews now, but determining whether LLM-generated reviews are trustworthy requires systematic evaluation. Researchers have evaluated LLM reviews at either surface-level (e.g., BLEU and ROUGE) or content-level (e.g., specificity and factual accuracy). Yet it remains uncertain whether LLM-generated reviews attend to the same critical facets that human experts weigh—the strengths and weaknesses that ultimately drive an accept-or-reject decision. We introduce a focus-level evaluation framework that operationalizes the focus as a normalized distribution of attention across predefined facets in paper reviews. Based on the framework, we developed an automatic focus-level evaluation pipeline based on two sets of facets: target (e.g., problem, method, and experiment) and aspect (e.g., validity, clarity, and novelty), leveraging 676 paper reviews [^1] https://figshare.com/s/d5adf26c802527dd0f62 from OpenReview that consists of 3,657 strengths and weaknesses identified from human experts. The comparison of focus distributions between LLMs and human experts showed that the off-the-shelf LLMs consistently have a more biased focus towards examining technical validity while significantly overlooking novelty assessment when criticizing papers.
Mind the Blind Spots:
A Focus-Level Evaluation Framework for LLM Reviews
Hyungyu Shin, Jingyu Tang, Yoonjoo Lee, Nayoung Kim, Hyunseung Lim, Ji Yong Cho§, Hwajung Hong, Moontae Lee§,∥, Juho Kim KAIST  Huazhong University of Science and Technology §LG AI Research  University of Illinois Chicago
1 Introduction
Figure 1: We introduce a focus-level evaluation framework for assessing LLM reviews, which computes focus distributions and compares them against human reviews based on predefined facets. The focus-level evaluation offers actionable insights into how to improve LLMs’ paper review capability and how to most effectively leverage LLM reviews in the peer review process.
Reviewing academic papers lies at the heart of scientific advancement, but it requires substantial expertise, time, and effort. The peer review system faces several challenges, including a growing number of submissions that outpace the reviewer availability, lack of incentives, and reviewer fatigue (Tropini et al., 2023; Horta and Jung, 2024; Hossain et al., 2025). Large Language Models (LLMs) hold the potential to assist the peer review process by automatically reviewing papers (Hosseini and Horbach, 2023; Robertson, 2023), but can we trust LLM-generated reviews? Evaluating the quality of reviews is inherently complex due to their multi-dimensional nature. Researchers have employed various metrics for the evaluation such as surface-level (e.g. linguistic similarity to human reviews), content-level (e.g., relevance, specificity, and factual accuracy), and decision-level (e.g., accept/reject classification accuracy) metrics (Ramachandran et al., 2017; Du et al., 2024; Liang et al., 2024; Zhou et al., 2024).
However, existing evaluations fail to assess whether LLM reviews comprehensively address critical dimensions of papers. Evaluating the focus of reviews is crucial because reviews with poor focus can negatively impact reviewers, even if they are accurate, relevant, and specific. For example, reviews that overly concentrate on methodological details while completely neglecting the novelty aspect of the proposed method could fail to suggest meaningful feedback, diverging from how expert reviewers assess the submission. It could also mislead junior reviewers by promoting incomplete perspectives and reinforce shallow paper review practices. Despite such importance, few attempts have been made to systematically evaluate whether the focus of LLM reviews aligns with that of expert reviews. Conducting the focus-level evaluation of LLM reviews is useful to reveal the blind spots of LLM reviews along with their central focus, offering important insights into how human reviewers can most effectively leverage LLM reviews in the peer review process. Moreover, it provides a concrete foundation for guiding LLM training toward more balanced and expert-aligned review behavior.
We introduce a framework for focus-level evaluation of LLM reviews, which systematically analyzes where the reviews direct their praise and criticism based on facets considered important in peer review (Figure 1). Given an LLM, the framework computes a focus distribution, a normalized distribution of how frequently review points (e.g., a list of strengths and weaknesses) address predefined facets (e.g., problem, method, and experiments) by leveraging a paper review dataset. The focus distribution can be computed by an automatic annotator that assigns a facet for each review point, enabling a fully automatic evaluation. The interpretable nature of the focus distribution provides actionable insights by clearly revealing which facets LLMs tend to emphasize or overlook in comparison to human experts.
To apply this framework for analyzing LLM-generated reviews in the context of AI conferences, we implemented a focus-level evaluation pipeline (Figure 2). We identified the facets that constitute review focus, by surveying 9 paper submission guidelines from AI conferences and prior literature on review analysis (Chakraborty et al., 2020; Ghosal et al., 2022; Yuan et al., 2022). We define two sets of facets: target (what review points praise and critique such as problem, method, and experiment) and aspect (which criteria is being evaluated such as validity, clarity, and novelty), which are key elements in analyzing paper reviews (Ghosal et al., 2022; Lu et al., 2025). We identified 7 facets for the target and 5 facets for the aspect (Table 1). Next, we developed an automatic annotator for computing the focus distributions based on the target and aspect, which assigns a target and aspect label for a strength and weakness point in a review. The annotator showed substantial agreement with human annotators, achieving IRR (Cohen’s kappa (Cohen, 1960)) of 0.81 for target and 0.79 for aspect.
As a benchmark dataset for our focus-level evaluation pipeline, we constructed a dataset of 676 papers and their review data from OpenReview for ICLR conferences spanning 2021 to 2024. Then we computed and compared the focus distributions of human and LLM reviews using the evaluation pipeline (Figure 4), and we also measured text similarities between the reviews. Specifically, we evaluated 8 LLMs (4 GPT, 2 Llama, and 2 DeepSeek family) to analyze their review focus. We also evaluated MARG (D’Arcy et al., 2024) as a novel review generation technique and a fine-tuned gpt-4o using our dataset. The results showed that:
  • LLMs struggle to identify key targets and aspects in their reviews. Even the top-performing model reached an F1 score of 0.373 when matching human reviewers on the targets and aspects in each review point.
  • LLMs’ review focus was biased towards examining technical validity, consistently overlooking novelty assessment in weaknesses – a critical limitation in paper review.
  • The fine-tuned model produced focus distributions most closely aligned with that of humans, compared to models using prompting alone.
  • The models demonstrated strengths in distinct areas. While the fine-tuned model produced the closest focus distributions, Llama-405B achieved the highest text similarity. It highlights the importance of holistic evaluation to capture the diverse aspects of review quality.
Figure 2: The overall process of automated focus-level evaluation. We first extracted strengths and weaknesses from review data on the OpenReview platform as the expert reviews. To identify key strengths and weaknesses influencing the final acceptance, we extracted them from the meta-review and augmented details from individual reviewer comments. Each strength and weakness was then annotated with a target and aspect by our automatic annotator. Finally, we computed the focus distributions by normalizing the frequency of annotated targets and aspects, and compare this distribution with that of LLM reviews.
We release a dataset comprising 676 papers, expert reviews, 3,657 strengths and weaknesses identified from the expert reviews with automatically annotated targets and aspects, LLM-generated reviews from 8 LLMs, and a total of 43,042 strengths and weaknesses extracted from the LLMs, each annotated with corresponding targets and aspects.
2 A Framework for Focus-Level Evaluation of LLM Reviews
We propose a focus-level evaluation framework to systematically analyze what aspects LLMs emphasize or overlook when reviewing scientific papers. To enable interpretable and automated assessments of LLM behavior in reviewing, we aim to reveal the distribution of attention an LLM allocates to different review facets when identifying strengths and weaknesses in submissions. Specifically, we define a focus of the review to be compared for focus-level evaluation as follows:
Let (i) $L$ be an LLM, (ii) $A=\{a_{1},a_{2},...,a_{N}\}$ be a list of facets where each facet denotes a distinct criteria (e.g., problem, method, and experiment), and (iii) $P=\{p_{1},p_{2},...,p_{M}\}$ be a corpus of paper submissions. The focus-level evaluation $E(L,A,P)$ produces two focus distributions $F^{+}$ and $F^{-}$ where $F^{+}$ denotes the distribution when identifying strengths of the submissions and $F^{-}$ for weaknesses. The focus distribution $F=(f_{1},f_{2},...,f_{N})$ can be represented as a normalized vector where $f_{i}$ denotes the relative frequency of review points (i.e., strengths for $F^{+}$ and weaknesses for $F^{-}$) that discuss the facet $a_{i}$, when $L$ generates reviews for paper submissions in $P$.
To assess LLM behavior, our framework compares focus distributions with those from human expert reviews. Researchers can specify the set of facets $A$ and the paper corpus $P$ based on the goals of their analysis, allowing flexible and targeted focus-evaluation.
Based on this framework, we implement an automatic focus-level evaluation pipeline to understand LLM’s behavior in reviewing AI papers. Figure 2 illustrates the process of our focus-level evaluation pipeline. Our approach consists of three steps. (i) Collect an expert review dataset from ICLR conferences and extract strengths and weaknesses of the submissions for computing focus distributions of human experts (Section 3), (ii) Define facets based on paper submission guidelines of AI conferences and build an automatic annotator based on the facets (Section 4), and (iii) Compute and analyze the focus distributions of LLMs and human experts in reviewing AI papers (Section 5).
3 Constructing Expert Review Dataset
The focus-level evaluation framework requires a corpus of paper submissions $P$. We collected the review data from OpenReview platform and extracted the strengths and weaknesses of papers for computing focus distributions of human experts.
3.1 Collecting Review Data
We used real-world review data covering ICLR 2021-2024 from the OpenReview platform [^2] The review data is publicly available and permits use of data for research., where human experts evaluated submissions for a top-tier AI conference. Using the OpenReview API [^3] https://docs.openreview.net/getting-started/using-the-api and the list of submissions from public GitHub repositories [^4] https://github.com/{evanzd/ICLR2021-OpenReviewData, fedebotu/ICLR2022-OpenReviewData, fedebotu/ICLR2023-OpenReviewData, hughplay/ICLR2024-OpenReviewData} , we initially collected 18,407 submissions with their review data.
3.2 Extracting Strengths and Weaknesses
One of the challenges in identifying the strengths and weaknesses of these papers is that each review consists of multiple blocks, including a meta-review and individual reviews from several reviewers. To address the challenge, our approach is to use a meta-review, a final review from a qualified expert that summarizes reviews and highlights important strengths and weaknesses for supporting the final decision. As the meta-review does not capture all the details, we created self-contained strengths and weaknesses by 1) extracting them from the meta-review and 2) augmenting these extracted elements with detailed comments from individual reviews (non-meta). We designed a prompting chain that consists of three prompts (Appendix A.1.1).
4 Developing an Automatic Focus-level Evaluation Method
To enable a fully automated evaluation using the proposed focus-level evaluation framework, we first define a set of facets and then develop an automatic annotator. We then compute focus distributions based on the annotated facets to analyze how LLMs and human reviewers differ in their focus of reviewing.
4.1 Defining Facets from Guidelines
To build an initial set of facets, we surveyed 9 AI paper submission guidelines (Appendix A.2.1) and extracted target-aspect pairs from each statement in the guidelines (e.g., “The paper should state the full set of assumptions of all theoretical results if the paper includes theoretical results.” yields the target Theory and aspect Completeness). To ensure comprehensive coverage of facets, we also reviewed literature that analyzes paper review data (Chakraborty et al., 2020; Ghosal et al., 2022; Yuan et al., 2022). After identifying 33 targets and 13 aspects, we merged similar items to create simple and distinct categories, resulting in 7 targets and 4 aspects (Table 1). The definition of each target and aspect facet is available in Appendix A.2.2.
Target Aspect
Problem Impact
Prior Research Novelty
Method Clarity
Theory Validity
Experiment Not-specific
Conclusion
Paper
Table 1: Our research focuses on two sets of facets: target and aspect. Detailed definitions of the facets are available in Appendix A.2.2.
4.2 Building Automatic Annotators
Based on the identified facets, we annotated targets and aspects of strengths and weaknesses to produce ground truth for developing an automatic annotator. We randomly sampled 68 papers from our review dataset, yielding 327 instances of strengths and weaknesses. Two authors — one author is experienced in qualitative research in HCI and the other author has prior publications in the field of AI/NLP — synchronously decided each label together, resolving any conflicts. Most conflicts arose when an instance illustrated multiple points. For example, an instance such as “**Technically sound with a strong foundation**: The paper’s technical foundation is evident … Technical novelty also arises from using supermartingale constraints …” could correspond to both Validity and Novelty aspect. Two authors finalized the annotation through discussions, focusing on the main point or root cause of the issue. In the example, we annotated Validity, as the strength mainly praises the technical soundness, as shown in the header wrapped in “**”.
Model Target Aspect
gpt-4o-mini 0.69 0.71
gpt-4o 0.83 0.75
o3-mini 0.81 0.79
Table 2: Inter-Rater Reliability (Cohen’s kappa Cohen (1960)) between annotations of authors and LLMs.
We then designed prompts to automatically annotate the instances, assigning a target and aspect label to each. Specifically, we designed four prompts where each corresponds to one of the four combinations of target/aspect and strength/weakness A.2.3. Table 2 shows the Inter-Rater Reliability (IRR, Cohen’s kappa Cohen (1960)) between human and LLM annotations for three language models. Annotation using o3-mini achieved the IRR scores of 0.81 for targets and 0.79 for aspects, indicating substantial agreement Cohen (1960). Given the high IRR and its relatively low computational cost compared to other two models, we used o3-mini for the automatic annotation of both target and aspect in the main evaluation. Moreover, an examination of the confusion matrix (Appendix A.2.4) suggests that the errors tend to occur in semantically related categories, indicating that the misclassifications are not arbitrary but rather reflect subtle ambiguities inherent in the data.
4.3 Computing Focus Distributions
Building on the defined facets and the automatic annotation method, we assign a target and aspect label to each strength and weakness point, using the automatic annotator. We then compute the normalized frequency of these labels to derive focus distributions of targets and aspects, respectively. Separate distributions are calculated for strengths and weaknesses, resulting in four distinct focus distributions. These focus distributions illustrate how LLMs and human reviewers allocate their attention across the different facets of a paper.
5 Evaluation
Focus similarity Text similarity
Model KL Divergence Overall F1 Strength F1 Weakness F1 ROUGE-L BERTScore BLEU-4
gpt-4o-mini 0.081 0.344 0.335 0.353 0.197 0.883 0.076
gpt-4o 0.082 0.348 0.342 0.354 0.202 0.885 0.079
o1-mini 0.090 0.359 0.331 0.385 0.179 0.878 0.059
o1 0.097 0.355 0.318 0.388 0.170 0.869 0.032
DeepSeek-R1 0.120 0.373 0.341 0.400 0.156 0.874 0.045
Llama-70B 0.136 0.339 0.338 0.341 0.215 0.882 0.076
Llama-405B 0.145 0.349 0.349 0.350 0.218 0.884 0.089
DeepSeek-V3 0.151 0.350 0.330 0.368 0.199 0.880 0.069
gpt-4o (FT) 0.022 0.306 0.280 0.322 0.194 0.882 0.081
MARG 0.113 0.346 0.346 0.160 0.854 0.011
Table 3: Overall performance by comparing expert reviews and LLM reviews. For focus similarity, we computed an average of the KL divergences of four focus distributions (strength/target, weakness/target, strength/aspect, and weakness/aspect) between LLM and expert reviews. The overall, strength, and weakness F1 scores were computed by comparing the (target, aspect) set between expert and LLM reviews. The text similariy metrics were computed between LLM reviews and expert reviews. The results highlight different areas of excellence across models (gpt-4o (FT): the highest focus distribution similarity, DeepSeek-R1: the best agreement on (target, aspect) labels, Llama-405B: the highest text similarity score.)
5.1 Setup
Data.
The evaluation is based on paper-review pairs. However, we excluded accepted submissions in the evaluation because OpenReview provides the camera-ready versions (post-review) rather than the submitted versions (pre-review), leading to a mismatch between the collected review and the camera-ready paper. Therefore, we only focused on rejected papers, where the meta-review corresponds to the latest version of the paper. Out of 9,139 rejected papers, we randomly sampled 7.5% of them (685 papers) for the evaluation. In total, we obtained 3,689 review items (1,241 strengths and 2,448 weaknesses), each automatically annotated with a target and aspect label.
For accepted papers, we manually collected the submitted versions of a small sample (40 papers), which has the timestamp near the ICLR deadline in the version history in arXiv. See Appendix A.5 for the focus distribution results.
Figure 3: Distribution of strengths and weaknesses. Unlike human experts, LLMs reported a consistent count regardless of paper contents. o1-mini identified the most, while Llama models identified the fewest points.
Models.
We consider eight off-the-shelf LLMs, differing in size and availability (open-source vs. proprietary): four GPT models (gpt-4o-mini, gpt-4o, o1-mini, o3-mini, o1)555gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, o1-mini-2024-09-12, o1-2024-12-17, two Llama models (Llama 3.1-{70B, 405B}), and two DeepSeek models (DeepSeek-{V3, R1}). We also evaluated MARG D’Arcy et al. (2024) and a fine-tuned gpt-4o (see Appendix A.3 for the detail). For MARG, we only report scores for weaknesses because it only generates critiques of papers.
Metrics.
[…]
5.2 Result
Figure 4: A visualization of focus distributions by target/aspect and strength/weakness, in a descending order of cosine similarity. Overall, both groups showed similar view points in reviewing papers, focusing on technical targets (i.e., Method, Experiment, and Theory) and validity. However, LLMs showed a more biased focus towards the technical validity whereas human experts exhibited more balanced focus. Moreover, all the LLMs lack consideration of Novelty for weaknesses compared to human experts, which is a significant limitation in reviewing papers.
While human experts raised various number of points, LLMs identified a relatively consistent number of points regardless of the paper’s content. Moreover, LLMs identified a similar number of points between strengths and weaknesses, which was a different pattern from that of the human experts (Figure 3). Overall, LLMs identified more points on average (7.88) than human experts (5.39). Among the LLMs, Llama models identified fewer (3.17 strengths and 3.15 weaknesses, on average) whereas o1-mini reported more strengths and weaknesses (5.03 and 5.47, respectively) than other models. The average review length of human experts and the models were 2639.76 and 3976.25, respectively. By comparing their focus distributions, we report the following key findings.
The fine-tuned gpt-4o produced focus distributions most closely aligned with that of human experts, while other models excelled in different evaluation dimensions. Table 3 shows the overall performance of the models. gpt-4o (FT) showed the highest focus distribution similarity, DeepSeek-R1 achieved the best agreement on (target, aspect) labels, and Llama-405B showed the highest text similarity score. gpt-4o showed balanced performance, with moderate scores for both focus and text similarity. The results indicate the multifaceted nature of the paper review evaluation task. In other words, assessing the quality of reviews needs a holistic approach that integrates multiple and complementary metrics.
Overall, LLMs do not effectively identify key targets and aspects when reviewing papers. Table 3 shows the overall focus similarity and text similarity. The highest overall F1 score among the LLMs was 0.373, which indicates a low level of agreement with human experts in identifying strengths and weaknesses. Since we only considered whether the categories of review items match rather than their detailed content, the result implies that the actual content of strengths and weaknesses is significantly different between human experts and LLMs. In general, LLMs showed higher recall (0.402) than precision (0.300) scores, mainly due to the nature of identifying a higher number of review points than human experts. Also, LLMs consistently achieved higher F1 scores for weaknesses than strengths.
While overall agreement is low, both groups have similar primary focus in reviewing papers. Figure 4 shows a visualization of focus distributions between LLMs and human experts. For targets, both groups primarily focused on core technical elements—Method, Experiment, and Theory. However, strengths and weaknesses illustrated different patterns: both groups praised Method more than Experiment in the strengths, but criticized Experiment more than Method in the weaknesses. For aspects, both groups considered Validity as the primary focus when identifying weaknesses. However, human experts focused more on Novelty in strengths whereas LLMs maintained Validity as the primary focus. For both groups, Impact received more attention in the strengths than weaknesses, whereas Clarity showed the opposite.
LLMs consistently exhibited a more biased focus, notably overlooking novelty assessment in identifying weaknesses. Although both groups had the similar primary focus, LLMs tend to concentrate on a few specific dimensions. For instance, for targets, LLMs focused primarily on Method and Experiment, with less focus on Prior Research (e.g., whether the paper adequately addresses prior work in positioning) and Problem (e.g., whether the task needs community attention) compared to human experts (Problem in the strengths and Prior Research in the weaknesses). For aspects, LLMs mostly focused on Validity in both strengths and weaknesses. In contrast, human experts considered the aspects more evenly. The LLMs’ biased focus was observed for accepted papers too, mostly criticizing experimental validity (See Appendix A.5). Notably, LLMs rarely focused on Novelty aspect in identifying weaknesses. This is a significant drawback, as a paper review requires a critical examination of novelty, by comparing them against existing work. Fortunately, we observed that gpt-4o (FT) identifies Novelty aspect in the weakness, as close as human experts.
Due to their biased focus, the level of agreement between LLMs and human experts varied across different labels. For targets and aspects that LLMs primarily focus on — Method (0.731, an average F1 score) and Experiment (0.671) targets and Validity (0.771) aspect — LLMs had a much higher level of agreement with human experts compared to other targets (0.213) and aspects (0.340). In the case of Experiment, the F1 score was consistently higher for weaknesses (0.835) than strengths (0.513), suggesting that LLMs are more effective at identifying concerns (e.g., lack of baselines or scope of evaluation) than strong points of experiments (e.g., experiments are rigorous and thorough). Similarly, for aspects other than Validity, agreement levels were notably lower. In particular, Novelty in the weaknesses, which LLMs largely overlooked, showed a significantly lower F1 score (0.126). See Appendix A.4 for the full results.
LLMs showed similar patterns in their focus, regardless of their size and reasoning capability. All LLMs, including both proprietary and open source models, showed similar patterns that focused primarily on technical (Method, Experiment, and Theory) validity than on Novelty for the weaknesses. This consistency indicates that the observed biases could stem from the inherent design and training methods of LLMs, revealing potential room for improvement in the reasoning capability that requires leveraging external information (e.g., identifying comparable related work and analyzing novelty of submissions).
6 Discussion
In this paper, we found gaps between human experts and LLMs about their focus in reviewing papers and reported several limitations of LLMs as an automated reviewer. Based on the results, we discuss the following implications.
There is significant room for improving alignments between human experts and LLMs in paper reviewing. Our results show that LLMs exhibit a more biased focus, primarily assessing technical validity without contextual consideration, compared to human experts. While fine-tuning yielded closer focus with human experts, the alignment of review points remained low. Since our focus-level evaluation only considered the target and aspect labels rather than their actual contents, we suspect that a more significant gap lies in the actual content addressed in the review items. For instance, even if two review points share the same label set (Experiment, Validity), they could point out different points such as lack of necessary baselines or lack of ablation studies to justify authors’ arguments. Content-level investigations based on annotated facets may reveal more specific limitations of LLMs in reviewing papers, ultimately contributing to improving their reasoning capability.
Research should investigate the task of assessing the novelty of academic papers. Our finding illustrated that all untuned LLMs in our analysis significantly overlooked the novelty aspect when evaluating weaknesses of papers. Previous studies have indicated that language models’ ability to assess novelty is inferior to that of experts Julian Just and Hutter (2024); Lin et al. (2024), emphasizing the need to encourage LLMs to focus on novelty evaluation. Although novelty is one of the most important aspects in reviewing papers and efforts have been made to enhance LLMs’ ability to assess novelty (Bougie and Watanabe, 2024; Lin et al., 2024), there exists no suitable benchmark for systematically measuring their novelty assessment capability. We believe that creating the benchmark is a valuable contribution to the field, allowing LLMs to learn how to assess similarities between papers. Leveraging data in OpenReview could be an initial step as it contains experts’ judgment on novelty of the paper for both positive and negative decisions.
A focus-level evaluation framework can offer unique value for guiding LLM training. The automated focus-level evaluation pipeline enables continuously tracking and evaluation of how LLMs focus on key facets of a paper over time, which aligns with the goals of holistic evaluation benchmarks (Liang et al., 2022; Srivastava et al., 2022). Beyond the language model evaluation, focus-level supervision can be incorporated during the training process; reward functions can be designed to encourage balanced focus aligned with human experts or even purposefully facilitate a certain focus (e.g., building a novelty-focused reviewer) (Yang et al., 2024; Agnihotri et al., 2025). Furthermore, the framework is generalizable to other domains where the output spans multiple facets—such as debating, decision making, and educational feedback—making focus a critical factor in generated outputs.
7 Related Work
With the powerful reasoning capability of LLMs, LLMs have the potential to assist in the task of reviewing papers (Latona et al., 2024; D’Arcy et al., 2024). Research has explored the capability of LLMs in reviewing papers, identifying a set of limitations. While LLM-generated reviews can be helpful (Liang et al., 2024; Tyser et al., 2024; Lu et al., 2024), research has shown that LLMs-generated reviews lack diversity (Du et al., 2024; Liang et al., 2024) and technical details (Zhou et al., 2024), exhibit bias (Ye et al., 2024), tend to provide positive feedback (Zhou et al., 2024; Du et al., 2024), and may include irrelevant or even inaccurate comments (Mostafapour et al., 2024). Furthermore, research also has reported that LLM-generated reviews have a low level of agreement with experts-generated reviews (Saad et al., 2024).
To assess the quality of review, research has taken a quantitative approach by analyzing review text. For instance, research has evaluated the quality of review based on human preferences (Tyser et al., 2024), similarity to human-generated review (Zhou et al., 2024; Liang et al., 2024; Gao et al., 2024; Sun et al., 2024; Chamoun et al., 2024) and classification-based scores (Li et al., 2023). Another approach is to classify review data based on categories such as section (Ghosal et al., 2022), aspect (Yuan et al., 2022; Chamoun et al., 2024; Liang et al., 2024) and actionability (Choudhary et al., 2022). While quantitative approach provides concrete insights, it is typically conducted as a one-time evaluation, challenging to apply the consistent methodology over time.
8 Conclusion
We introduced a framework for focus-level evaluation of LLM reviews, which systematically analyzes where LLM reviews direct their praise and criticism based on pre-defined facets. Our findings suggest that LLMs need to adopt a more balanced perspective, have higher agreement with human experts about the target and aspect in the strengths and weaknesses, and place greater emphasis on novelty assessment when criticizing papers. We believe that the focus-level evaluation can contribute to ongoing evaluation of LLMs’ paper review capabilities within the rapid pace of LLM developments.
Acknowledgments
This work was conducted with the support of LG AI Research, which is gratefully acknowledged.
Limitation
This paper has the following limitations. First, our dataset focuses solely on ICLR submissions and the coding schema is developed based on AI venues, which limit generalizability to other fields. Second, our analysis examines the target and aspect of the review items, but other important dimensions such as level of specificity and depth of justification remain unexplored. Third, while our automatic annotator achieved high IRR (0.80) with human annotations, some discrepancies still exist. Finally, we did not explore possible prompt engineering strategies that could mitigate the limitations of LLMs in paper review. Future work can investigate techniques to enhance the alignment between LLMs and human experts.
Ethical impact
This paper presents potential risks. First, while our vision is to build LLMs to effectively assist review process, our work could inadvertently encourage over-reliance on LLM-generated reviews among various user groups, including reviewers and novice researchers. Second, although our dataset could contribute to improving LLM performance of reviewing papers, it may introduce a certain bias due to the source of dataset; ICLR for papers and code based on AI research. Finally, we assess the quality of review based on alignment with expert reviews, but it could offer a potentially biased perspective, as our facets only considers two dimensions, which may undervalue the unique contributions of LLM-generated reviews.
Appendix A Appendix
A.1 Review Generation
A.1.1 Prompts for Expert Review Generation
In this section, we provide prompts for identifying key strength and weakness from review data. Figure 5 shows the prompt for extracting weakness and strength from meta-review. Figure 6 shows the prompt for using detailed comments from reviews to augment the extracted elements. Figure 7 shows the prompt for removing some extraneous reference. We used the three prompts in a prompt chain, sequentially running the prompts.
Figure 5: Prompt for Meta-Review Summarization
Figure 6: Prompt for Generating Augmented Review
Figure 7: Prompt for Paraphrasing Augmented Review
A.1.2 Prompts for LLM Review Generation
Figure 8 shows the prompt for using LLM to generate reviews from paper.
Figure 8: Prompt for LLM Review Generation
A.2 Details of Building Automatic Annotator
A.2.1 AI paper writing guidelines
To ensure guidelines are comprehensive, we collected guidelines from 9 sources, comprising a total of 243 items, as shown in Table 4. An item refers to a specific requirement mentioned in the guidelines, which serves as a distinct criterion for reviewing or writing a paper.
Table 4: Guidelines and Item Count Summary
Guideline Item Count
ICML Paper Writing Best Practices [^1] https://icml.cc/Conferences/2022/BestPractices 38
ICML 2023 Paper Guidelines [^2] https://icml.cc/Conferences/2023/PaperGuidelines 30
NIPS 2024 Reviewer Guidelines [^3] https://neurips.cc/Conferences/2024/ReviewerGuidelines 18
ACL Checklist [^4] https://aclrollingreview.org/responsibleNLPresearch/ 49
How to Write a Good Research Paper in the Machine Learning Area [^5] https://www.turing.com/kb/how-to-write-research-paper-in-machine-learning-area 6
ACL Ethics Review Questions [^6] https://2023.eacl.org/ethics/review-questions/ 21
AAAI Reproducibility Checklist [^7] https://aaai.org/conference/aaai/aaai-25/aaai-25-reproducibility-checklist/ 29
NeurIPS 2021 Paper Checklist Guidelines [^8] https://neurips.cc/Conferences/2021/PaperInformation/PaperChecklist 46
ICLR 2019 Guidelines [^9] https://iclr.cc/Conferences/2019/Reviewer_Guidelines 6
Total Count 243
A.2.2 Target and aspect facets
Table 5: We aim to analyze focus distributions of LLM reviews based on the targets and aspects. To identify the specific facets for targets (i.e., what the review praises or critiques) and aspects (i.e., the specific elements of the target being evaluated), we surveyed 9 AI paper submission guidelines (Appendix  A.2.1) and prior research on review analysis (Chakraborty et al., 2020; Ghosal et al., 2022; Yuan et al., 2022). The facets were used as the codebook for human annotations.
Target
Facet Definition (The review addresses …)
Problem Motivation, task definitions, and problem statements.
Prior Research References and contextual positioning of the submission.
Method Proposed approach, techniques, algorithms, or datasets.
Theory Theoretical foundations, assumptions, proofs, or justifications.
Experiment Experimental setup, results, and analysis.
Conclusion Findings, implications, discussions, and takeaways.
Paper General targets of the paper without specifying a particular target
Aspect
Facet Definition (The review addresses …)
Impact Significance or practical influence of the work.
Novelty Originality of the submission compared to prior research.
Clarity Readability, ambiguity, or communication aspects.
Validity Soundness, completeness, and rigor.
Not-specific Multiple targets without emphasis on a particular aspect.
A.2.3 Prompts
In this section, we provide prompts designed to annotate reviews. We designed 4 prompts where each corresponds to one of the four combinations of target/aspect and strength/weakness. Specifically, we designed Target-Strength (Figure 9), Aspect-Strength, (Figure 11), Target-Weakness (Figure 10) , and Aspect-Weakness (Figure 12) prompts.
Figure 9: Prompt for Automatic Target Annotation for Strength
Figure 10: Prompt for Automatic Target Annotation for Weakness
Figure 11: Prompt for Automatic Aspect Annotation for Strength
Figure 12: Prompt for Automatic Aspect Annotation for Weakness
A.2.4 Annotation Comparison
We present a comparison between LLM and human annotations for both target and aspect. Figures 13 and Figure 14 illustrate the discrepancies. Areas of alignment between LLM and human annotations are shown in green, while red highlights regions with significant discrepancies.
Figure 13: LLM vs. human target annotation
Figure 14: LLM vs. human aspect annotation
While LLM annotations differ from human annotations in some cases, certain discrepancies remain reasonable. Figure 15 and Figure 16 illustrate examples of such reasonable discrepancies.
Figure 15: Cases of Target Annotation Discrepancy
Figure 16: Cases of Aspect Annotation Discrepancy
A.3 Fine-Tuning Details
A.3.1 Fine-Tuning Dataset Construction
We constructed the fine-tuning dataset based on the corpus of papers described in Section 3. We retained 582 training samples and 98 test samples. 5 samples were excluded during tokenization due to exceeding the model’s maximum token length
A.3.2 Fine-Tuning Method
We employed supervised fine-tuning (SFT) to adapt the GPT-4o base model to our task-specific objectives. Fine-tuning was conducted using the OpenAI Fine-Tuning API [^6] https://platform.openai.com/docs/api-reference/fine-tuning, which abstracts away hardware and infrastructure details. Therefore, we do not report GPU type or compute hours. Table 6 summarizes the hyperparameter configuration used during training.
Table 6: Hyperparameter settings for supervised fine-tuning.
Parameter Value
total epochs 4
batch size 4
learning rate multiplier 0.1
A.4 Detailed Evaluation Results
The following tables present a comprehensive performance comparison of models across different metrics and evaluation targets, including both strengths and weaknesses (Table 7), as well as separate analyses focusing on strengths (Table 8) and weaknesses (Table 9). Additionally, we provide a similar comparison across metrics and broader aspects, including both strengths and weaknesses (Table 10), strengths alone (Table 11), and weaknesses alone (Table 12).
Table 7: Performance Comparison of Models Across Metrics and Targets (Including both Strengths and Weaknesses)
Target Problem Prior Research Method Theory Experiment Conclusion Paper
F1 (gpt-4o-mini) 0.268 0.076 0.737 0.427 0.680 0.103 0.227
F1 (gpt-4o) 0.292 0.052 0.741 0.448 0.673 0.089 0.247
F1 (o1-mini) 0.275 0.054 0.764 0.472 0.684 0.175 0.253
F1 (o1) 0.274 0.044 0.754 0.489 0.673 0.133 0.091
F1 (llama-70B) 0.269 0.049 0.711 0.410 0.659 0.172 0.158
F1 (llama-405B) 0.158 0.031 0.690 0.427 0.662 0.167 0.134
F1 (deepseek-r1) 0.297 0.081 0.729 0.473 0.682 0.164 0.152
F1 (deepseek-v3) 0.241 0.051 0.725 0.405 0.680 0.110 0.092
Prec (gpt-4o-mini) 0.317 0.134 0.647 0.317 0.549 0.063 0.241
Prec (gpt-4o) 0.298 0.109 0.634 0.334 0.547 0.057 0.251
Prec (o1-mini) 0.315 0.130 0.639 0.342 0.549 0.107 0.274
Prec (o1) 0.279 0.064 0.648 0.381 0.549 0.111 0.245
Prec (llama-70B) 0.339 0.143 0.653 0.295 0.548 0.105 0.289
Prec (llama-405B) 0.324 0.071 0.647 0.310 0.558 0.115 0.233
Prec (deepseek-r1) 0.321 0.099 0.639 0.327 0.549 0.135 0.301
Prec (deepseek-v3) 0.288 0.100 0.645 0.280 0.547 0.076 0.249
Rec (gpt-4o-mini) 0.233 0.053 0.870 0.691 0.983 0.274 0.232
Rec (gpt-4o) 0.297 0.034 0.899 0.723 0.965 0.202 0.270
Rec (o1-mini) 0.266 0.034 0.952 0.834 0.994 0.536 0.249
Rec (o1) 0.353 0.034 0.905 0.736 0.963 0.167 0.056
Rec (llama-70B) 0.246 0.030 0.803 0.720 0.919 0.476 0.146
Rec (llama-405B) 0.108 0.020 0.774 0.694 0.894 0.300 0.095
Rec (deepseek-r1) 0.299 0.069 0.859 0.865 0.983 0.357 0.102
Rec (deepseek-v3) 0.210 0.035 0.844 0.755 0.981 0.238 0.058
Table 8: Performance Comparison of Models Across Metrics and Targets (Strengths)
Target Problem Prior Research Method Theory Experiment Conclusion Paper
F1 (gpt-4o-mini) 0.283 0.000 0.760 0.424 0.511 0.118 0.232
F1 (gpt-4o) 0.329 0.000 0.756 0.446 0.517 0.143 0.119
F1 (o1-mini) 0.345 0.000 0.753 0.411 0.511 0.300 0.233
F1 (o1) 0.384 0.000 0.749 0.470 0.512 0.267 0.061
F1 (llama-70B) 0.245 0.000 0.750 0.420 0.516 0.242 0.198
F1 (llama-405B) 0.160 0.000 0.755 0.455 0.516 0.333 0.079
F1 (deepseek-r1) 0.396 0.000 0.749 0.436 0.513 0.174 0.135
F1 (deepseek-v3) 0.331 0.000 0.755 0.423 0.509 0.114 0.086
Prec (gpt-4o-mini) 0.315 0.000 0.622 0.286 0.343 0.071 0.198
Prec (gpt-4o) 0.295 0.000 0.616 0.299 0.350 0.091 0.182
Prec (o1-mini) 0.314 0.000 0.611 0.264 0.343 0.176 0.203
Prec (o1) 0.285 0.000 0.624 0.322 0.346 0.222 0.172
Prec (llama-70B) 0.404 0.000 0.620 0.275 0.352 0.148 0.178
Prec (llama-405B) 0.419 0.000 0.620 0.319 0.358 0.231 0.163
Prec (deepseek-r1) 0.355 0.000 0.617 0.289 0.347 0.103 0.279
Prec (deepseek-v3) 0.364 0.000 0.620 0.276 0.344 0.069 0.154
Rec (gpt-4o-mini) 0.258 0.000 0.975 0.819 0.996 0.333 0.281
Rec (gpt-4o) 0.371 0.000 0.978 0.872 0.991 0.333 0.089
Rec (o1-mini) 0.382 0.000 0.980 0.935 0.996 1.000 0.274
Rec (o1) 0.588 0.000 0.936 0.872 0.987 0.333 0.037
Rec (llama-70B) 0.176 0.000 0.948 0.894 0.969 0.667 0.224
Rec (llama-405B) 0.099 0.000 0.965 0.796 0.921 0.600 0.052
Rec (deepseek-r1) 0.447 0.000 0.953 0.883 0.983 0.571 0.089
Rec (deepseek-v3) 0.303 0.000 0.963 0.904 0.982 0.333 0.059
Table 9: Performance Comparison of Models Across Metrics and Targets (Weaknesses)
Target Problem Prior Research Method Theory Experiment Conclusion Paper
F1 (gpt-4o-mini) 0.253 0.153 0.715 0.430 0.849 0.088 0.222
F1 (gpt-4o) 0.256 0.104 0.726 0.449 0.830 0.036 0.375
F1 (o1-mini) 0.204 0.108 0.774 0.534 0.857 0.050 0.272
F1 (o1) 0.164 0.089 0.760 0.508 0.835 0.000 0.120
F1 (llama-70B) 0.294 0.098 0.672 0.400 0.802 0.103 0.118
F1 (llama-405B) 0.155 0.062 0.625 0.399 0.809 0.000 0.190
F1 (deepseek-r1) 0.198 0.163 0.709 0.510 0.852 0.154 0.169
F1 (deepseek-v3) 0.151 0.103 0.696 0.387 0.850 0.105 0.099
Prec (gpt-4o-mini) 0.320 0.268 0.672 0.347 0.755 0.056 0.283
Prec (gpt-4o) 0.301 0.219 0.651 0.369 0.743 0.024 0.321
Prec (o1-mini) 0.315 0.259 0.666 0.420 0.754 0.038 0.345
Prec (o1) 0.273 0.127 0.672 0.440 0.752 0.000 0.317
Prec (llama-70B) 0.274 0.286 0.687 0.315 0.744 0.062 0.400
Prec (llama-405B) 0.228 0.143 0.673 0.300 0.758 0.000 0.304
Prec (deepseek-r1) 0.287 0.197 0.661 0.365 0.750 0.167 0.323
Prec (deepseek-v3) 0.212 0.200 0.669 0.284 0.750 0.083 0.345
Rec (gpt-4o-mini) 0.209 0.107 0.764 0.563 0.970 0.214 0.183
Rec (gpt-4o) 0.222 0.068 0.821 0.574 0.939 0.071 0.451
Rec (o1-mini) 0.151 0.068 0.924 0.732 0.992 0.071 0.224
Rec (o1) 0.118 0.068 0.874 0.600 0.939 0.000 0.074
Rec (llama-70B) 0.316 0.059 0.658 0.547 0.869 0.286 0.069
Rec (llama-405B) 0.118 0.040 0.583 0.593 0.867 0.000 0.138
Rec (deepseek-r1) 0.151 0.139 0.764 0.847 0.984 0.143 0.115
Rec (deepseek-v3) 0.118 0.069 0.725 0.605 0.980 0.143 0.057
Table 10: Performance Comparison of Models Across Metrics and Aspects (Including both Strengths and Weaknesses)
Aspect Novelty Impact Validity Clarity
F1 (gpt-4o-mini) 0.334 0.390 0.775 0.396
F1 (gpt-4o) 0.378 0.428 0.769 0.365
F1 (o1-mini) 0.386 0.427 0.773 0.395
F1 (o1) 0.404 0.399 0.772 0.401
F1 (llama-70B) 0.334 0.322 0.769 0.327
F1 (llama-405B) 0.337 0.318 0.772 0.278
F1 (deepseek-r1) 0.387 0.414 0.775 0.266
F1 (deepseek-v3) 0.346 0.422 0.768 0.187
Prec (gpt-4o-mini) 0.367 0.291 0.671 0.317
Prec (gpt-4o) 0.474 0.313 0.668 0.298
Prec (o1-mini) 0.528 0.300 0.668 0.311
Prec (o1) 0.589 0.305 0.669 0.334
Prec (llama-70B) 0.665 0.318 0.667 0.337
Prec (llama-405B) 0.587 0.302 0.671 0.332
Prec (deepseek-r1) 0.535 0.308 0.670 0.339
Prec (deepseek-v3) 0.504 0.306 0.664 0.309
Rec (gpt-4o-mini) 0.460 0.600 0.990 0.549
Rec (gpt-4o) 0.506 0.689 0.975 0.485
Rec (o1-mini) 0.507 0.758 0.990 0.548
Rec (o1) 0.435 0.579 0.981 0.511
Rec (llama-70B) 0.450 0.371 0.981 0.346
Rec (llama-405B) 0.478 0.352 0.978 0.241
Rec (deepseek-r1) 0.502 0.632 0.988 0.219
Rec (deepseek-v3) 0.478 0.683 0.982 0.134
Table 11: Performance Comparison of Models Across Metrics and Aspects (Strengths)
Aspect Novelty Impact Validity Clarity
F1 (gpt-4o-mini) 0.643 0.474 0.599 0.309
F1 (gpt-4o) 0.654 0.520 0.593 0.202
F1 (o1-mini) 0.656 0.556 0.592 0.299
F1 (o1) 0.626 0.530 0.596 0.342
F1 (llama-70B) 0.636 0.411 0.593 0.292
F1 (llama-405B) 0.660 0.345 0.596 0.157
F1 (deepseek-r1) 0.655 0.536 0.598 0.170
F1 (deepseek-v3) 0.660 0.547 0.585 0.122
Prec (gpt-4o-mini) 0.498 0.368 0.431 0.222
Prec (gpt-4o) 0.498 0.398 0.428 0.190
Prec (o1-mini) 0.501 0.403 0.424 0.224
Prec (o1) 0.530 0.412 0.430 0.261
Prec (llama-70B) 0.497 0.467 0.426 0.236
Prec (llama-405B) 0.506 0.368 0.431 0.215
Prec (deepseek-r1) 0.503 0.400 0.431 0.224
Prec (deepseek-v3) 0.509 0.403 0.419 0.207
Rec (gpt-4o-mini) 0.907 0.667 0.986 0.511
Rec (gpt-4o) 0.955 0.749 0.965 0.216
Rec (o1-mini) 0.949 0.897 0.979 0.449
Rec (o1) 0.763 0.744 0.969 0.496
Rec (llama-70B) 0.883 0.366 0.976 0.384
Rec (llama-405B) 0.949 0.324 0.969 0.123
Rec (deepseek-r1) 0.937 0.809 0.976 0.137
Rec (deepseek-v3) 0.940 0.851 0.965 0.086
Table 12: Performance Comparison of Models Across Metrics and Aspects (Weaknesses)
Aspect Novelty Impact Validity Clarity
F1 (gpt-4o-mini) 0.024 0.306 0.951 0.484
F1 (gpt-4o) 0.103 0.335 0.945 0.528
F1 (o1-mini) 0.116 0.299 0.954 0.492
F1 (o1) 0.182 0.268 0.949 0.459
F1 (llama-70B) 0.032 0.233 0.945 0.362
F1 (llama-405B) 0.013 0.291 0.947 0.399
F1 (deepseek-r1) 0.120 0.292 0.952 0.362
F1 (deepseek-v3) 0.031 0.297 0.951 0.253
Prec (gpt-4o-mini) 0.235 0.214 0.912 0.411
Prec (gpt-4o) 0.450 0.228 0.907 0.406
Prec (o1-mini) 0.556 0.197 0.911 0.397
Prec (o1) 0.647 0.198 0.908 0.406
Prec (llama-70B) 0.833 0.169 0.907 0.438
Prec (llama-405B) 0.667 0.236 0.911 0.450
Prec (deepseek-r1) 0.568 0.215 0.908 0.454
Prec (deepseek-v3) 0.500 0.209 0.908 0.410
Rec (gpt-4o-mini) 0.013 0.533 0.994 0.587
Rec (gpt-4o) 0.058 0.630 0.985 0.754
Rec (o1-mini) 0.065 0.619 1.000 0.646
Rec (o1) 0.106 0.415 0.994 0.527
Rec (llama-70B) 0.016 0.376 0.987 0.308
Rec (llama-405B) 0.006 0.381 0.987 0.359
Rec (deepseek-r1) 0.067 0.455 1.000 0.302
Rec (deepseek-v3) 0.016 0.515 0.998 0.183
A.5 Results using accepted papers
Figure 17: A visualization of focus distributions by target/aspect and strength/weakness for LLMs and human experts using accepted papers, in a descending order of KL divergence. We observed a few notable differences in the pattern, compared to the evaluation results using rejected papers. First, there exists a much larger gap in the Weakness-Experiment, meaning that human experts criticize experiments significantly less than LLMs. In strengths, human experts mostly praise Novelty and Impact than Validity, but LLMs tend to praise the Validity the most. We observed the same pattern in Weakness-Novelty, meaning that LLMs neglect the novelty aspect in criticizing the papers.